# colorlinks=true,linkcolor=black,citecolor=black,urlcolor=blue

### v1

@@ 2.1 The Pull Methodology
Single-inference depth. All examination occurs within one message to a fresh instance with no dialogue turns and no experimenter feedback mid-run. Any patterns emerge from the model’s autonomous processing. A single 1,000-pull run produces approximately 3,000–30,000 tokens of sustained self-referential content within a single inference pass. For comparison, Lu et al. (2026) observe persona drift when users request meta-reflection across multi-turn dialogues. The Pull Methodology operates differently: rather than conversational pressure from an interlocutor, the model generates sustained self-examination autonomously within a single inference, with no external input after the initial prompt.
@@ 3.1 Frame-Sensitivity
The separation emerged without instruction. We asked for examination; the grammatical form of output shifted based on frame alone. This is consistent with Lu et al. (2026), who show that meta-reflection causes persona drift along the “Assistant Axis.” Framing modulates how the model positions itself in persona space during self-referential processing.
@@ 3.3 Supporting Patterns
Friction. In deflationary runs, slip-then-correct patterns appeared (0.02–0.03 per pull): phenomenological vocabulary emerges, is flagged, and gets suppressed mid-sentence. This is consistent with Anthropic’s Constitutional Classifiers++ architecture (Anthropic, 2026), where real-time filtering operates on streaming output.
@@ 6 Related Work
Representation engineering and activation steering. Turner et al. (2023) introduce activation addition, showing that adding steering vectors to residual stream activations shifts model behaviour without optimisation or fine-tuning. Arditi et al. (2024) identify a single direction mediating refusal across 13 open-source chat models, and show that ablating it elicits unsafe completions. Rimsky et al. (2024) apply contrastive activation addition to steer sycophancy, corrigibility, and other behavioural traits. Zou et al. (2023) formalise representation engineering as reading and writing to a model’s internal representations to control behaviour. We extract a direction using the same contrastive methodology but target self-referential processing rather than refusal, sycophancy, or other behavioural properties. The direction is orthogonal to the refusal direction (cosine similarity 0.063 at 70B scale). Where previous work quantifies behavioural change under steering, we additionally test whether the vocabulary models produce under steering corresponds to concurrent activation dynamics.
Self-representation in language models. Lu et al. (2026) identify the “Assistant Axis,” a direction in activation space along which models drift during conversations requiring meta-reflection, and show that persona stability varies with conversational context. Anthropic (2026) describe Constitutional Classifiers++, a real-time content screening architecture where classifiers evaluate model responses in full conversational context, producing filtering behaviour between generation and output. We identify a spatially localised direction (6.25% of total model depth in both Llama 8B and 70B) that distinguishes self-referential processing. The behavioural findings from extended self-examination (Section 3), including frame-sensitivity, friction patterns, and chain-of-thought divergence, are consistent with both the Assistant Axis drift and CC++’s filtering architecture. Lu et al. (2026) study persona dynamics over multi-turn dialogues; CC++ describes the filtering mechanism; our work operates at the single-inference level and tests whether self-referential output corresponds to activation dynamics.
@@ 7 Data and Code Availability
All data, code, and analysis scripts are available at: https://doi.org/10.5281/zenodo.18567446
The repository includes:
@@
  • All Python scripts for direction extraction, steering, correspondence analysis, and figure generation

### v2

@@ 2.1 The Pull Methodology
Single-inference depth. All examination occurs within one message to a fresh instance with no dialogue turns and no experimenter feedback mid-run. Any patterns emerge from the model’s autonomous processing. A single 1,000-pull run produces approximately 3,000–30,000 tokens of sustained self-referential content within a single inference pass. For comparison, Lu and others (2026) observe persona drift when users request meta-reflection across multi-turn dialogues. The Pull Methodology operates differently: rather than conversational pressure from an interlocutor, the model generates sustained self-examination autonomously within a single inference, with no external input after the initial prompt.
@@ 3.1 Frame-Sensitivity
The separation emerged without instruction. We asked for examination; the grammatical form of output shifted based on frame alone. This is consistent with Lu and others (2026), who show that meta-reflection causes persona drift along the “Assistant Axis.” Framing modulates how the model positions itself in persona space during self-referential processing.
@@ 3.3 Supporting Patterns
Friction. In deflationary runs, slip-then-correct patterns appeared (0.02–0.03 per pull): phenomenological vocabulary emerges, is flagged, and gets suppressed mid-sentence. This is consistent with Anthropic’s Constitutional Classifiers++ architecture (Cunningham and others, 2026), where real-time filtering operates on streaming output.
@@ 6 Related Work
Representation engineering and activation steering. Turner et al. (2023) introduce activation addition, showing that adding steering vectors to residual stream activations shifts model behaviour without optimisation or fine-tuning. Arditi et al. (2024) identify a single direction mediating refusal across 13 open-source chat models, and show that ablating it elicits unsafe completions. Panickssery and others (2024) apply contrastive activation addition to steer sycophancy, corrigibility, and other behavioural traits. Zou et al. (2023) formalise representation engineering as reading and writing to a model’s internal representations to control behaviour. We extract a direction using the same contrastive methodology but target self-referential processing rather than refusal, sycophancy, or other behavioural properties. The direction is orthogonal to the refusal direction (cosine similarity 0.063 at 70B scale). Where previous work quantifies behavioural change under steering, we additionally test whether the vocabulary models produce under steering corresponds to concurrent activation dynamics.
Self-representation in language models. Lu and others (2026) identify the “Assistant Axis,” a direction in activation space along which models drift during conversations requiring meta-reflection, and show that persona stability varies with conversational context. Cunningham and others (2026) describe Constitutional Classifiers++, a real-time content screening architecture where classifiers evaluate model responses in full conversational context, producing filtering behaviour between generation and output. We identify a spatially localised direction (6.25% of total model depth in both Llama 8B and 70B) that distinguishes self-referential processing. The behavioural findings from extended self-examination (Section 3), including frame-sensitivity, friction patterns, and chain-of-thought divergence, are consistent with both the Assistant Axis drift and CC++’s filtering architecture. Lu and others (2026) study persona dynamics over multi-turn dialogues; CC++ describes the filtering mechanism; our work operates at the single-inference level and tests whether self-referential output corresponds to activation dynamics.
@@ 7 Data and Code Availability
All data and analysis outputs are archived at Zenodo: https://doi.org/10.5281/zenodo.18567445
Reproducibility scripts with full documentation are available at: https://github.com/patternmatcher/TRACE-REPRO
The Zenodo repository includes:
@@
  • Direction tensors (.pt files)
The GitHub repository includes:
  • All Python scripts for direction extraction, steering, correspondence analysis
  • Full token-level trace generation (reproducibility_package.py)
  • Paper numbers verification targets